OpenStudy (katwqyz):

Rewrite the equation in slope-intercept form. (y=mx+b) -y + 4x = - 8

OpenStudy (5sos):

So first what you would do is you would subtract 4x from both sides. When you do that what do you get?

KjSaif (kjsaif):

^^

OpenStudy (katwqyz):

I HAVE NO IDEA.

OpenStudy (katwqyz):

Please walk me thru.

OpenStudy (katwqyz):

I never listened to my teacher now i'm in a hole#

OpenStudy (5sos):

Ok so if you did that you would get -y=-4x-8 But now you need to get rid of the negative in front of the y. So you multiply the whole equation by -1 and that would give you your final answer. y-4x+8

OpenStudy (5sos):

*y=4x+8

RhondaSommer (rhondasommer):

\[-y + 4x = - 8\]subrtact 4x on both sides of the equal sign.\[-y+4x-4x=-4x-8\\-y=-4x-8\]divide -1 on both sides of the equal sign\[\frac{-y}{-1}=\frac{-4x-8}{-1}\\y=4x+8\] thats your answer! :)

OpenStudy (katwqyz):

ok I think I get it

OpenStudy (katwqyz):

Thanks to both of you.

OpenStudy (katwqyz):

uh can you guys like give each other a medal